5-2x=3x-7x+25
5(d+4)=7(d-2)
3g-3=3(7-g)



I really just dont understand equation with both sides. So explain CLEARLY please.

Answer by stanbon(75887)

5-2x=3x-7x+25
Combine like terms on the same side:
5-2x = -4x+25
Add 4x to both sides:
2x+5 = 25
Subtract 5 from both sides
2x = 20
Divide both sides by 2
x = 10
=================

5(d+4)=7(d-2)
Used the distributive law
5d+20 = 7d-14
Subtract 5d from both sides
20 = 2d-14
Add 14 to both sides
2d = 34
Divide both sides by 2
d = 17
==================
3g-3=3(7-g)
Divide both sides by 3 to get:
g - 1 = 7-g
Add g to both sides
2g - 1 = 7
Add 1 to both sides
2g = 8
Divide both sides by 2
g = 4
=============
Cheers,
Stan H.